A train travelling along a straight track starts from rest at point A and accelerates uniformly to 20m/s in 20s. It travels at this speed for 60s then slows down uniformly to rest in 40s at point C. it stays at rest at C for 30s, then reverses direction accelerating uniformly to 10m/s in 10s. It travels at this speed for 30s, then slows down uniformly to rest in 10s when it reaches point B.
a) Plot a graph of the motion of the train
b) Use your graph to calculate
i) The trains displacement from point A when it reaches point C
ii) The trains displacement from point A when it reaches point B
iii) The trains acceleration each time its speed changes
1
Expert's answer
2019-10-23T09:50:20-0400


Let's find the train's displacement from point A when it reaches point C. We should calculate the area of the trapeze ADEC:

"S_1=\frac{60+120}{2}\cdot{}20=1800" m.

Now let's find the train's displacement from point A when it reaches point B. We should calculate the difference of areas of the trapeze ADEC and trapeze FGHB:

"S_2=S_1-\frac{50+30}{2}\cdot{}10=1800-400=1400" m.

Let's find the trains acceleration each time its speed changes.

"0\eqslantless{}t\eqslantless{}20," "a_1=\frac{20-0}{20}=1" m/s2.

"80\eqslantless{}t\eqslantless{}120," "a_2=\frac{0-20}{40}=-0.5" m/s2.

"150\eqslantless{}t\eqslantless{}160," "a_3=\frac{-10-0}{10}=-1" m/s2.

"190\eqslantless{}t\eqslantless{}200," "a_4=\frac{0-(-10)}{10}=1" m/s2.